# Forced 3D reconnection in an exponentially separating magnetic field

### v1

@@ Abstract
We present a solvable scenario for 3D reconnection in a sheared magnetic field. We consider a localized external force that is applied slowly and then maintained, generating an Alfvénic perturbation that spreads along the field lines. Separation of the sheared field lines causes the scale of the perturbation across the field to decrease, enabling magnetic diffusion to be enhanced. For a fusion-motivated equilibrium with exponential field-line separation, we find a reconnection timescale proportional to $\mathcal{S}/\ln\mathcal{S}$ under magnetohydrodynamics (MHD) and to $\mathcal{S}^{1/3}$ for semi-collisional electron-only reconnection, where $\mathcal{S}$ is the Lundquist number of the perturbed flux tube. We generalize these results to arbitrary magnetic geometries, showing that the latter is geometry-independent. Interestingly, we find that slower field-line separation yields an increased reconnection rate in MHD.
@@
Introduction.  The rate of magnetic reconnection—i.e., of topological reconfiguration of a magnetic field in a conducting medium—in a fully three-dimensional geometry is an issue of great significance both for fusion devices and for magnetized astrophysical systems. In this Letter, we present a solvable scenario for the 3D resistive reconnection of separating field lines. We consider a force applied locally to a narrow, sheared flux tube, inducing a resistively damped Alfvén-wave packet that thins as it propagates. Because the packet is associated with a change to the local magnetic-field direction, its resistive diffusion modifies the connectivity of the global field.
@@
We focus primarily on a particular fusion-motivated equilibrium [Eq. (2)] with field lines that separate exponentially with distance [Fig. 1(a)]. It has been suggested, and is perhaps intuitively compelling, that the rate of reconnection of exponentially separating field lines is fast, i.e., at most logarithmic in the Lundquist number $\mathcal{S}$ (the ratio of the ideal to nominal resistive timescales) [1, 2]. Indeed, naïvely, it seems intuitive that reconnection induced by a diffusive Alfvén wave, as described above, would be fast in this case: the wave would be sheared to resistive scales in $\ln\mathcal{S}$ Alfvén timescales. However, we find that the reconnection timescale remains proportional to a finite power of $\mathcal{S}$ (see abstract). At the end of the Letter, we generalize to other geometries, finding faster reconnection for field lines that separate more slowly.
@@
Figure 1: Panel (a): The magnetic geometry considered in this study [Eq. (2)]. The surface of constant $\alpha$ along which we consider displacing a field line is shown with $\psi$ increasing from blue to red. Arrows show the projection of the magnetic-field direction onto the $xy$-plane. Panel (b): Projection onto the $xy$-plane of the field line through $(-\sqrt{2}a,0,0)$ at (i) $\tilde{t}=0$ (dots with black outlines) and (ii) $\tilde{t}\gg\tau_{\mathrm{rec}}$ (blue-red line). Note that the orientation of the $y$-axis in panel (b) is different than in panel (a).
@@
Clebsch coordinates. We represent the equilibrium magnetic field $\bm{B}_{0}$ with Clebsch coordinates $\psi$ and $\alpha$, i.e., $\bm{B}_{0}=\bm{\nabla}\psi\times\bm{\nabla}\alpha$. We take $\psi$ to have units of magnetic flux and $\alpha$ to be dimensionless. The Clebsch coordinates are constant along unperturbed field lines and so can be used to label them; the arc length $\ell$ along the line completes the coordinate set. The distance perpendicular to ${\bm{B}}_{0}$ between the field lines $(\psi,\alpha)$ and $(\psi+d\psi,\alpha)$ is
$\mathrm{d}l_{\perp}=\frac{|\bm{\nabla}\alpha|}{B_{0}}\mathrm{d}\psi.$ (1)
The rate of separation of field lines on a surface of constant $\alpha$ (“$\alpha$ surface”) is therefore encoded in the dependence of $|\bm{\nabla}\alpha|$ on $\ell$. This dependence will set the rate of magnetic reconnection in our calculation.
@@
The equilibrium (2) is visualized in Fig. 1(a). Toroidicity notwithstanding, each lobe resembles the magnetic field of a tokamak in a divertor configuration (with the lobes taken together, it resembles a “doublet” [3]). We expect the results of our Letter to be applicable to reconnection in such a device, forced, for example, by an ELM-control coil (whose effect on the topology of edge fields and transport may be crucial for sustained confinement [4, 5, 6]).
We consider specifically reconnection in the vicinity of the field line passing through the point $(-\sqrt{2}a,0,0)$ on the separatrix. As we show in Section 1 of the Supplementary Information, we may define $\alpha$ surfaces such that
@@
$\ell_{0}=a^{2}B_{\mathrm{ext}}/4I_{0}$ is the magnetic-shear length (we take $\ell_{0}=a$ in all numerical results presented in this Letter) and $B_{0}(\ell)=|\bm{B}_{0}|$. According to Eqs. (1) and (3), field lines separate exponentially with $\ell$ as ${\ell\to\infty}$.
Dynamical model. We take the equilibrium magnetic field to be strong, and therefore consider a perturbation that is extended along the field line. We consider two dynamical regimes: the MHD case, in which the perpendicular scale of the perturbation is much larger than the ion gyroradius $\rho_{i}$, and the kinetic case, in which it is much smaller than $\rho_{i}$. The dynamical equations that we employ are those of low-$\beta$ reduced MHD [7] in the MHD case and the semi-collisional limit of the Kinetic Reduced Electron Heating Model (KREHM) of Ref. [8] in the kinetic case. The latter regime is interesting a priori because the parallel group velocity of kinetic Alfvén waves is an increasing function of their perpendicular wavenumber [9]: the equilibrium shear causes them to accelerate and thus reach diffusive scales more quickly  [^1] We note that the semi-collisional limit (collisionless ions but collisional electrons) means that these waves are immune to Landau damping..
We assume an eikonal form for the velocity streamfunction $\Phi$ and perturbed magnetic-flux function $\Psi$, viz., $\Phi(\psi,\alpha,\ell,t)=\bar{\Phi}(\ell,t)e^{in\alpha}$ and $\Psi(\psi,\alpha,\ell,t)=\bar{\Psi}(\ell,t)e^{in\alpha}$, where $n\gg 1$ (cf. Refs. [11, 12, 13, 14]). As we explain in Section 2 of the Supplementary Information, this choice yields
@@
The variables that appear in Eqs. (5) and (6) are as follows. The velocity in the $\alpha$ surface is ${\delta\bm{u}_{\perp}=\bm{b}_{0}\times\bm{\nabla}\Phi=u(\psi,\alpha,\ell,t)e^{in\alpha}\bm{b}_{0}\times\bm{\nabla}\alpha}$; ${\delta\bm{B}_{\perp}=\bm{b}_{0}\times\bm{\nabla}\Psi=\delta B_{\perp}(\psi,\alpha,\ell,t)B_{0}(\ell)e^{in\alpha}\bm{b}_{0}\times\bm{\nabla}\alpha}$ is the magnetic-field perturbation in the same surface; $\bm{b}_{0}(\ell)=\bm{B}_{0}/B_{0}$; and $J_{\parallel}=|\bm{\nabla}\alpha|^{2}\delta B_{\perp}$ is the perturbed parallel current. Tildes signify normalization to the dimensionally appropriate combination of the shear length $\ell_{0}$ and the ideal timescale $\tau_{A}$. The latter is the shear-length crossing time of, in the MHD case, an Alfvén wave (i.e., $\tau_{A}=\ell_{0}/v_{A,\mathrm{ext}}$, with $v_{A,\mathrm{ext}}$ the Alfvén speed associated with $B_{\mathrm{ext}}$) or, in the kinetic case, a kinetic Alfvén wave (i.e., $\tau_{A}=\sqrt{2}\ell_{0}^{2}/n\rho_{i}v_{A,\mathrm{ext}}$). Similarly normalized, $\tilde{F}(\tilde{\ell},\tilde{t})$ represents a force with the same eikonal dependence on $\alpha$ as $\Phi$ and $\Psi$, which pushes the field line along the $\alpha$ surface. We take $\tilde{F}$ to have odd parity in $\tilde{\ell}$ and to turn on smoothly on the (normalized) timescale $\epsilon^{-1}\gg 1$, reaching a constant value at $\tilde{t}\gg\epsilon^{-1}$. In the numerical results shown in Figs. 1(b), 2 and 3, we choose $\tilde{B}_{0}(\tilde{\ell})^{2}\tilde{F}(\tilde{\ell},\tilde{t})=\tilde{\ell}^{3}e^{-\tilde{\ell}^{2}}(1-e^{-\epsilon^{2}\tilde{t}^{2}})$, absorbing the amplitude of $\tilde{F}$ into the definitions of $\tilde{u}$, $\delta\tilde{B}_{\perp}$ and $\tilde{J}_{\parallel}$. The Lundquist number for the perturbed flux tube, $\mathcal{\mathcal{S}}$, is
@@
where $S=v_{A,\mathrm{ext}}\ell_{0}/\eta$ is the Lundquist number of the global equilibrium, $\eta$ is the magnetic diffusivity and ${\rho_{\star}\equiv\rho_{i}/\ell_{0}}$ is the normalized ion gyroradius. The numerical values quoted in Eq. (7) correspond to $\mathcal{S}\sim 10^{8}$ and $\rho_{\star}\sim 10^{-2}$, which are typical for a tokamak.
@@
Reconnected flux.  Eq. (6) admits an invariant in the “ideal” limit $\mathcal{S}\to\infty$, given by
@@
$\Delta\psi$ is the difference (“step”) in the coordinate $\psi$ of the perturbed field line between $\ell=-\infty$ and $\ell=+\infty$: $\Delta\psi\neq 0$ implies reconnection of field lines from inside the separatrix with those outside it [see Fig. 1(b)]. In a fusion device, this would cause the loss of confinement of hot plasma from inside the separatrix by streaming. Our goal is to determine the evolution of $\Delta\psi$ for finite $\mathcal{S}\gg 1$.
@@
Late-time diffusive solutions. We consider the late-time ($\epsilon\tilde{t}\gg 1$) evolution of a diffusively spreading pulse induced by the force $\tilde{F}$. We plot numerical solutions of Eqs. (5) and (6) in Fig. 2. We now determine these solutions and their reconnection rates [Fig. 3] analytically.
We decompose the solutions into two regions: an inner region, where $\ell\sim\ell_{0}$ and the fluid inertia [left-hand side of Eq. (5)] can be neglected (because $\epsilon\ll 1$), and an outer region, where diffusion [last term in Eq. (6)] dominates over induction [first term in Eq. (6)].
@@
Figure 2: Evolution of $\tilde{u}$ and $\tilde{J}_{\parallel}$ for $\tilde{t}\sim\tau_{\mathrm{rec}}$. The panels on the left show the MHD case, those on the right show the kinetic case. Solid lines show numerical solutions of Eqs. (5) and (6) obtained with a finite-difference scheme. Dots show the theoretical prediction derived in the main text; dots with black outlines correspond to the outer solution [Eqs. (19) and (20) for the MHD case (corrected for the resistive diffusion of the inner solution, as described in Section 4 of the Supplementary information), Eqs. (25) and (26) for the kinetic case], dots with white outlines correspond to the inner solution [Eqs. (9) and (10)].
@@
MHD case: (i) Inner region (ideal). Neglecting inertia, Eq. (5) yields
$\tilde{J}_{\parallel\mathrm{in}}(\tilde{\ell},\tilde{t})=-\int_{-\infty}^{\tilde{\ell}}\mathrm{d}\tilde{\ell}^{\prime}\tilde{F}+\tilde{J}_{\parallel 0}(\tilde{t}),$ (9)
where $\tilde{J}_{\parallel 0}(\tilde{t})$ is the current at ~$\tilde{\ell}\gg 1$. From (6) we have
@@
$\tilde{u}_{\mathrm{in}}=\frac{1}{\tilde{B}_{0}(\tilde{\ell)}}\int_{0}^{\tilde{\ell}}\mathrm{d}\tilde{\ell}^{\prime}\frac{\tilde{B}_{0}(\tilde{\ell}^{\prime})}{|\tilde{\bm{\nabla}}\alpha|^{2}}\left(\frac{\partial\tilde{J}_{\parallel\mathrm{in}}}{\partial\tilde{t}}+\frac{\tilde{J}_{\parallel\mathrm{in}}}{\mathcal{S}}|\tilde{\bm{\nabla}}\alpha|^{2}\right).$ (10)
@@
The inner solution therefore depends on a single unknown function of time, $J_{\parallel 0}(t)$, which we shall determine by matching to the outer solution. We can deduce the value of $J_{\parallel 0}(t)$ at early times by assuming that the forcing-ramp-up time $\epsilon^{-1}$ is small compared with the normalized reconnection timescale $\tau_{\mathrm{rec}}$, whatever the latter turns out to be (we show how to relax this restriction in Section 3 of the Supplementary Information). In that case, $\Delta\psi$ is unchanged between $\tilde{t}=0$ and $\epsilon^{-1}\ll\tilde{t}\ll\tau_{\mathrm{rec}}$. At such times, Eqs. (8) and (9) require that
@@
$\tilde{J}_{\parallel 0}(t)=\mathcal{J}_{0}\equiv|\Delta^{\prime}|\int_{-\infty}^{\infty}\mathrm{d}\tilde{\ell}\frac{\tilde{B}_{0}(\tilde{\ell})}{|\tilde{\bm{\nabla}}\alpha|^{2}}\int_{-\infty}^{\tilde{\ell}}\mathrm{d}\tilde{\ell}^{\prime}\lim_{\tilde{t}\to\infty}\tilde{F},$ (11)
@@
MHD case: (ii) Outer region (diffusive).  At ~$\tilde{\ell}\gg 1$, the force $\tilde{F}$ is zero. Because $\epsilon\ll 1$, we also neglect the induction term in Eq. (6) compared with the diffusive one. In this case, $\tilde{u}_{\mathrm{out}}$ satisfies a diffusion equation:
$\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{t}}=\frac{\mathcal{S}}{|\tilde{\bm{\nabla}}\alpha|^{2}}\frac{\partial^{2}\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}^{2}},\quad|\tilde{\bm{\nabla}}\alpha|^{2}=\Lambda e^{2\tilde{\ell}},$ (12)
where $\Lambda=\exp[2\lim_{\tilde{\ell}\to\infty}(\tilde{\ell}(\tilde{z})-\tilde{z})]/4$ is a constant to be determined from Eq. (4). Eq. (12) admits a similarity variable $\xi={\Lambda}{e^{2\tilde{\ell}}}/{4\mathcal{S}\tilde{t}}$. Changing variables from $\tilde{\ell}$ to $\xi$, Eq. (12) becomes
@@
MHD case: (iii) Matching solutions, $\tilde{t}\ll\mathcal{S}$. For ${\tilde{t}\ll\mathcal{S}}$, resistive diffusion of the inner solution [the final term in Eq. (10)] can be neglected. The matching condition is then [from Eqs. (9), (10) and $\tilde{J}_{\parallel\mathrm{out}}=\mathcal{S}\partial\tilde{u}_{\mathrm{out}}/\partial\tilde{\ell}$]
$\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}=\lim_{\tilde{\ell}\to\infty}\tilde{u}_{\mathrm{in}}=\frac{\mathcal{S}}{|\Delta^{\prime}|}\frac{\mathrm{d}}{\mathrm{d}\tilde{t}}\lim_{\tilde{\ell}\to 0}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}}.$ (16)
Using standard results for Tricomi’s function, we have
$\lim_{\xi\to 0}\tilde{u}_{\mathrm{out}}=-\sum_{n}\frac{C_{n}\tilde{t}^{p_{n}}\ln(\Theta_{n}\xi)}{\Gamma(1+p_{n})},\quad\Theta_{n}\equiv e^{2\gamma_{\mathrm{E}}+\mathrm{F}\left(1+n\right)},$ (17)
up to $O(\xi)$, where $\gamma_{\mathrm{E}}$ is the Euler–Mascheroni constant and $\mathrm{F}$ is the digamma function. For $\tilde{t}\ll\mathcal{S}$, the logarithm in Eq. (17) is constant and approximately equal to $\ln(\Lambda/4\mathcal{S})$. Eqs. (16) and (17) together require that $p_{n+1}=p_{n}+1$, $C_{n+1}=-C_{n}/\tau_{\mathrm{rec}}$, where

### v2

@@ Abstract
We present a solvable scenario for 3D reconnection in a sheared magnetic field. We consider a localized external force that is applied slowly to a flux tube and then maintained, generating an Alfvénic perturbation that spreads along the field lines. Separation of the sheared field lines reduces the scale of the perturbation across the field, enhancing magnetic diffusion. For a fusion-motivated equilibrium with exponential field-line separation, we find a reconnection timescale proportional to $\mathcal{S}/\ln\mathcal{S}$ under magnetohydrodynamics (MHD) and to $\mathcal{S}^{1/3}$ for semicollisional electron-only reconnection, where $\mathcal{S}$ is the Lundquist number of the perturbed flux tube. We generalize these results to arbitrary magnetic geometries, showing that the semicollisional case is geometry independent. Interestingly, we find that slower field-line separation yields an increased reconnection rate in MHD.
@@
Introduction.  Magnetic reconnection—the topological reconfiguration of a magnetic field in a conducting medium—is a fundamental process of key importance in both laboratory and astrophysical plasmas. In fusion devices, reconnection can cause sawtooth crashes [1, 2] and act both as trigger [3] and saturation mechanism [4] for edge-localized modes (ELMs). In astrophysical and space plasmas, reconnection mediates and drives turbulence [5, 6], accelerates particles [7] and enables solar flares [8]. While reconnection in 2D is enabled by the tearing instability [9, 10] (particularly in its nonlinear [11] and/or kinetic [12] regimes), it has been suggested that alternative 3D mechanisms may facilitiate faster reconnection [13, 14, 15]. In this Letter, we present a solvable case of 3D reconnection occurring via a 3D analogue of resistive tearing. The process we consider occurs when a force is applied locally to a sheared flux tube and maintained. This induces a resistively damped Alfvén-wave packet that thins as it propagates along separating field lines. Because the packet is associated with a local change in the direction of the magnetic field, its resistive diffusion modifies the connectivity of the global field.
@@
Although we generalize to arbitrary equilibria at the end of the Letter, we develop and illustrate our results for a particular fusion-motivated equilibrium [Eq. (2)] with field lines that separate exponentially with the distance $\ell$ along them [Fig. 1(a)]. It has been suggested, and is perhaps intuitively compelling, that the rate of reconnection of exponentially separating field lines is fast, i.e., at most logarithmic in the Lundquist number $\mathcal{S}$ (the ratio of the ideal to nominal resistive timescales) [13, 14]. It seems intuitive that this would be the case for the scenario described above: a diffusing Alfvén wave would be sheared to resistive scales in $\ln\mathcal{S}$ Alfvén timescales. However, we find that the reconnection timescale remains proportional to a finite power of $\mathcal{S}$ (see abstract) because the field-line separation arrests propagation of the wave. Indeed, we find faster reconnection for field lines that separate more slowly (provided that they separate faster than $\ell^{1/2}$).
@@
Figure 1: Panel (a): The magnetic geometry considered in this study [Eq. (2)]. The $\alpha$ surface along which we consider displacing a field line is shown with $\psi$ increasing from red to blue. Arrows show the projection of the magnetic-field direction onto the $xy$-plane. Panel (b): Projection onto the $xy$-plane of the field line through $(-\sqrt{2}a,0,0)$ at (i) $\tilde{t}=0$ (dots with black outlines) and (ii) $\tilde{t}\gg\tau_{\mathrm{rec}}$ (purple-green line). Note that the orientation of the $y$-axis in panel (b) is opposite to that in panel (a).
@@
Clebsch coordinates. We represent the equilibrium magnetic field $\bm{B}_{0}$ with Clebsch coordinates $\psi$ and $\alpha$, i.e., $\bm{B}_{0}=\bm{\nabla}\psi\times\bm{\nabla}\alpha$. We take $\psi$ to have units of magnetic flux and $\alpha$ to be dimensionless. The Clebsch coordinates are constant along unperturbed field lines and can therefore be used to label them; the arc length $\ell$ along the field line completes the coordinate set. We take the applied force $\bm{f}$ to be localized to, and directed along, a particular $\alpha$-surface (i.e., $\bm{f}\propto\bm{B}_{0}\times\bm{\nabla}\alpha$). The distance perpendicular to ${\bm{B}}_{0}$ between two field lines $(\psi,\alpha)$ and $(\psi+d\psi,\alpha)$ in an $\alpha$ surface is
$\mathrm{d}l_{\perp}=\frac{|\bm{\nabla}\alpha|}{B_{0}}\mathrm{d}\psi,$ (1)
so the rate of separation of the field lines is encoded in the dependence of $|\bm{\nabla}\alpha|$ on $\ell$. This dependence determines the reconnection rate in our theory. (Note that the coordinate $\alpha$ is not uniquely defined for given $\bm{B}_{0}$. Different definitions correspond to different applied forces, which may be associated with different reconnection rates.)
@@
The equilibrium (2) is visualized in Fig. 1(a). Toroidicity notwithstanding, each lobe resembles the magnetic field of a tokamak in a divertor configuration (with the lobes taken together, it resembles a “doublet” [16]). We expect the results of our Letter to be applicable to reconnection in such a device, forced, for example, by an ELM-control coil (whose effect on the topology of edge fields and transport may be crucial for confinement [17, 18, 19]).
We consider an odd-parity force that acts in the vicinity of the point $(-\sqrt{2}a,0,0)$ on the separatrix, inducing reconnection of field lines passing through this region. As we explain in Section 1 of the End Matter, we may define $\alpha$ such that
@@
$\ell_{0}=a^{2}B_{\mathrm{ext}}/4I_{0}$ is the magnetic-shear length (we take $\ell_{0}=a$ in all numerical results presented in this Letter) and $B_{0}(\ell)=|\bm{B}_{0}|$. This choice corresponds to a force in the $\hat{\bm{x}}$ direction. According to Eqs. (1) and (3), field lines separate exponentially with $\ell$ as ${\ell\to\infty}$ [see Fig. 1(a)].
Dynamical model. We take the equilibrium magnetic field to be strong, and therefore consider a perturbation that is extended along the field line. We consider two dynamical regimes: the MHD case, in which the perpendicular scale of the perturbation is much larger than the ion gyroradius $\rho_{i}$, and the kinetic case, in which it is much smaller than $\rho_{i}$. We employ the equations of low-$\beta$ reduced MHD [20] to describe the MHD case and the semi-collisional limit of the Kinetic Reduced Electron Heating Model (KREHM) of Ref. [21] to describe the kinetic case (see Section 1 of the Supplementary Information). The latter regime is interesting a priori because the parallel group velocity of kinetic Alfvén waves is an increasing function of their perpendicular wavenumber [22]: the equilibrium shear causes them to accelerate and thus reach diffusive scales more quickly  [^1] We note that the semi-collisional limit (collisionless ions but collisional electrons) means that these waves are immune to Landau damping..
We assume an eikonal form for the velocity streamfunction $\Phi$ and perturbed magnetic-flux function $\Psi$, viz., $\Phi(\psi,\alpha,\ell,t)=\bar{\Phi}(\ell,t)e^{in\alpha}$ and $\Psi(\psi,\alpha,\ell,t)=\bar{\Psi}(\ell,t)e^{in\alpha}$, where $n\gg 1$ (cf. Refs. [24, 25, 26, 27]). As we explain in Section 1 of the Supplementary Information, this yields
@@
The variables that appear in Eqs. (5) and (6) are as follows. The velocity, directed parallel to the $\alpha$ surface, is ${\delta\bm{u}_{\perp}=\bm{b}_{0}\times\bm{\nabla}\Phi=u(\psi,\alpha,\ell,t)e^{in\alpha}\bm{b}_{0}\times\bm{\nabla}\alpha}$; ${\delta\bm{B}_{\perp}=\bm{b}_{0}\times\bm{\nabla}\Psi=\delta B_{\perp}(\psi,\alpha,\ell,t)B_{0}(\ell)e^{in\alpha}\bm{b}_{0}\times\bm{\nabla}\alpha}$ is the magnetic-field perturbation in the same surface; $\bm{b}_{0}(\ell)=\bm{B}_{0}/B_{0}$; and $J_{\parallel}=|\bm{\nabla}\alpha|^{2}\delta B_{\perp}$ is the perturbed parallel-current density. Tildes signify normalization to the dimensionally appropriate combination of the shear length $\ell_{0}$ and the ideal timescale $\tau_{A}$. The latter is the shear-length crossing time of, in the MHD case, an Alfvén wave (i.e., $\tau_{A}=\ell_{0}/v_{A,\mathrm{ext}}$, with $v_{A,\mathrm{ext}}$ the Alfvén speed associated with $B_{\mathrm{ext}}$) or, in the kinetic case, a kinetic Alfvén wave (i.e., $\tau_{A}=\sqrt{2}\ell_{0}^{2}/n\rho_{i}v_{A,\mathrm{ext}}$). Similarly normalized, $\tilde{F}(\tilde{\ell},\tilde{t})$ represents a force with the same eikonal dependence on $\alpha$ as $\Phi$ and $\Psi$, which pushes the field line along the $\alpha$ surface. We take $\tilde{F}$ to have odd parity in $\tilde{\ell}$ and to turn on smoothly on the (normalized) timescale $\epsilon^{-1}\gg 1$, reaching a constant value at $\tilde{t}\gg\epsilon^{-1}$. In the numerical results shown in Figs. 1(b), 2 and 3, we choose $\tilde{B}_{0}(\tilde{\ell})^{2}\tilde{F}(\tilde{\ell},\tilde{t})=\tilde{\ell}^{3}e^{-\tilde{\ell}^{2}}(1-e^{-\epsilon^{2}\tilde{t}^{2}})$, absorbing the amplitude of $\tilde{F}$ into the definitions of $\tilde{u}$, $\delta\tilde{B}_{\perp}$ and $\tilde{J}_{\parallel}$. We define the Lundquist number for the perturbed flux tube to be $\mathcal{\mathcal{S}}=\tau_{A}/\tau_{\eta}$, where $\tau_{\eta}=\ell_{0}^{2}/n^{2}\eta$ is the resistive timescale at the perpendicular scale $\ell_{0}/n$. Thus,
@@
where $S=v_{A,\mathrm{ext}}\ell_{0}/\eta$ is the Lundquist number of the global equilibrium, $\eta$ is the magnetic diffusivity and ${\rho_{\star}\equiv\rho_{i}/\ell_{0}}$ is the normalized ion gyroradius. The numerical values quoted in Eq. (7) correspond to $S\sim 10^{8}$ and $\rho_{\star}\sim 10^{-2}$, which are typical for a tokamak.
@@
Reconnected flux.  In the absence of the resistive final term, Eq. (6) conserves
@@
$\Delta\psi$ is the difference (“step”) in the coordinate $\psi$ of the perturbed field line between $\ell=-\infty$ and $\ell=+\infty$: $\Delta\psi\neq 0$ implies reconnection of field lines from inside the separatrix with those outside it [see Fig. 1(b)]. In a fusion device, this would cause the loss of confinement of hot plasma from inside the separatrix by streaming. The total flux reconnected through the separatrix [see Fig. 1(b)] is $\int\mathrm{d}\alpha\mathrm{d}\psi\sim\Delta\psi/n$. Our goal is to determine the evolution of $\Delta\psi$ for finite $\mathcal{S}\gg 1$.
@@
Late-time diffusive solutions. We consider the late-time ($\epsilon\tilde{t}\gg 1$) evolution of a diffusively spreading pulse induced by the force $\tilde{F}$. We plot numerical solutions of Eqs. (5) and (6) in Fig. 2. In what follows, we determine these solutions and their reconnection rates (Fig. 3) for $\mathcal{S}\gg 1$ analytically, using matched asymptotic expansions. There are two relevant solution domains: an inner region, ~$\tilde{\ell}\ll\ln\mathcal{S}$, where inertia [left-hand side of Eq. (5)] is negligible, and an outer region, ~$\tilde{\ell}\gg\ln(\ln\mathcal{S})$, where induction [first term in Eq. (6)] is. We justify neglecting these terms in the above-stated domains a posteriori—see Section 3 of the End Matter.
@@
Figure 2: Evolution of $\tilde{u}$ and $\tilde{J}_{\parallel}$ for $\tilde{t}\sim\tau_{\mathrm{rec}}$ in the MHD case. Solid lines show numerical solutions of Eqs. (5) and (6) obtained with a finite-difference scheme. Dots show the theoretical prediction derived in the main text: dots with black outlines correspond to the outer solution [Eqs. (19) and (20) (corrected for the resistive diffusion of the inner solution, as described in Section 4 of the End Matter)], dots with white outlines to the inner solution [Eqs. (9) and (10)].
@@
MHD case: (i) Inner region (~$\tilde{\ell}\ll\ln\mathcal{S}$). Neglecting inertia, Eq. (5) yields
$\tilde{J}_{\parallel\mathrm{in}}(\tilde{\ell},\tilde{t})=-\int_{-\infty}^{\tilde{\ell}}\mathrm{d}\tilde{\ell}^{\prime}\tilde{F}+\tilde{J}_{\parallel 0}(\tilde{t});$ (9)
substituting this into (6) yields, for $\epsilon\tilde{t}\gg 1$,
@@
$\tilde{u}_{\mathrm{in}}=\frac{1}{\tilde{B}_{0}}\int_{0}^{\tilde{\ell}}\mathrm{d}\tilde{\ell}^{\prime}\frac{\tilde{B}_{0}}{\mathcal{S}}\left(\frac{\mathcal{S}}{|\tilde{\bm{\nabla}}\alpha|^{2}}\frac{\mathrm{d}\tilde{J}_{\parallel 0}}{\mathrm{d}\tilde{t}}+\tilde{J}_{\parallel 0}-\int^{\ell^{\prime}}_{-\infty}\mathrm{d}\ell^{\prime\prime}\tilde{F}\right).$ (10)
@@
The inner solution depends on a single unknown function of time, $\tilde{J}_{\parallel 0}(t)$, which we shall determine by matching to the outer solution. We can deduce the value of $\tilde{J}_{\parallel 0}(\tilde{t})$ at early times by assuming that the forcing-ramp-up time $\epsilon^{-1}$ is small compared with the normalized reconnection timescale $\tau_{\mathrm{rec}}$, whatever the latter turns out to be (we show how to relax this restriction in Section 2 of the Supplementary Information). In that case, $\Delta\psi$ is unchanged between $\tilde{t}=0$ and $\epsilon^{-1}\ll\tilde{t}\ll\tau_{\mathrm{rec}}$. At such times, Eqs. (8) and (9) require that
@@
$\tilde{J}_{\parallel 0}(\tilde{t})=\mathcal{J}_{0}\equiv|\Delta^{\prime}|\int_{-\infty}^{\infty}\mathrm{d}\tilde{\ell}\frac{\tilde{B}_{0}(\tilde{\ell})}{|\tilde{\bm{\nabla}}\alpha|^{2}}\int_{-\infty}^{\tilde{\ell}}\mathrm{d}\tilde{\ell}^{\prime}\lim_{\tilde{t}\to\infty}\tilde{F},$ (11)
@@
Figure 3: Evolution of $\Delta\psi$ [Eq. (8)] as a function of time in the MHD (left) and kinetic (right) cases in the numerical solution of Eqs. (5) and (6). At $\tilde{t}\gg 1/\epsilon$, the evolution is as predicted by Eqs. (22) and (23) (see insets, which show time rescaled by the appropriate function of $\mathcal{S}$). Blue regions indicate $\tilde{t}\lesssim\epsilon^{-1}$, when the force $\tilde{F}$ is increasing in size. The power laws for $\Delta\psi$ in these regions are derived in Section 2 of the Supplementary Information.
MHD case: (ii) Outer region (~$\tilde{\ell}\gg\ln(\ln\mathcal{S})$). Dropping $\tilde{F}$ in Eq. (5) (as ~${\tilde{\ell}\gg 1}$) and neglecting induction in Eq. (6), $\tilde{u}_{\mathrm{out}}$ satisfies the diffusion equation
$\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{t}}=\frac{\mathcal{S}}{|\tilde{\bm{\nabla}}\alpha|^{2}}\frac{\partial^{2}\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}^{2}}.$ (12)
We now specialize to the equilibrium (2). For ~${\tilde{\ell}\gg 1}$, Eq. (3) reads |~${|\tilde{\bm{\nabla}}\alpha|^{2}=\Lambda e^{2\tilde{\ell}}}$, where ${\Lambda=\exp[2\lim_{\tilde{\ell}\to\infty}(\tilde{\ell}(\tilde{z})-\tilde{z})]/4}$ is obtained from Eq. (4). Eq. (12) then admits a similarity variable $\xi={\Lambda}{e^{2\tilde{\ell}}}/{4\mathcal{S}\tilde{t}}$. Changing variables from $\tilde{\ell}$ to $\xi$, Eq. (12) becomes
@@
MHD case: (iii) Matching solutions. Both the inner and outer solutions are valid for ${1\ll\ln(\ln\mathcal{S})\ll\tilde{\ell}\ll\ln\mathcal{S}}$ and must therefore match there. In this range, $\tilde{u}$ decreases linearly with $\tilde{\ell}$, i.e., ${\tilde{u}\to c_{1}(\tilde{t})-c_{2}(\tilde{t})\tilde{\ell}}$ (Fig. 2). Apart from at very late times $\tilde{t}\gtrsim\mathcal{S}$ [see “MHD case (v)”], only the first term in Eq. (10) for $\tilde{u}_{\mathrm{in}}$ contributes to $c_{1}$; the other two (resistive) terms determine $c_{2}$. From the outer solution, $c_{1}=\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}$. Equating the two expressions for $c_{1}$, the matching condition is
$\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}=\frac{\mathcal{S}}{|\Delta^{\prime}|}\frac{\mathrm{d}}{\mathrm{d}\tilde{t}}\lim_{\tilde{\ell}\to 0}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}},$ (16)
where we have used that $\tilde{J}_{\parallel 0}=\mathcal{S}\partial\tilde{u}_{\mathrm{out}}/\partial\tilde{\ell}$ in the matching region. Using the above-stated solutions of Eq. (15) and standard results for limits of Tricomi’s function, Eq. (14) implies that
$\lim_{\xi\to 0}\tilde{u}_{\mathrm{out}}=-\sum_{n}\frac{C_{n}\tilde{t}^{p_{n}}\ln(\Theta_{n}\xi)}{\Gamma(1+p_{n})}+O(\xi),$ (17)
where $\Theta_{n}\equiv e^{2\gamma_{\mathrm{E}}+\mathrm{F}\left(1+n\right)}$, $\gamma_{\mathrm{E}}$ is the Euler–Mascheroni constant and $\mathrm{F}$ is the digamma function. For $\tilde{t}\ll\mathcal{S}$, the logarithm in Eq. (17) is constant and approximately equal to $\ln(\Lambda/4\mathcal{S})$. Eqs. (16) and (17) together require that $p_{n+1}=p_{n}+1$, $C_{n+1}=-C_{n}/\tau_{\mathrm{rec}}$, where